# MAD-Sherlock: Multi-Agent Debates for Out-of-Context Misinformation Detection

### v1

@@ 3.3.1 API-Based Information Retrieval
The Bing Visual Search API (vis ) is used for the task of obtaining web pages related to a given image. A given image from the dataset is used to obtain a list of web pages completely and partially related to the image. We take the top three matching web pages in which the image appears. We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used. Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al. (2021)), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results. In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base. Since this is not the case for a significant fraction of examples in the dataset, it does not adversely affect system performance.
@@ 3.3.2 Summarization using LLM
Once the top three web pages have been identified, we scrape the text from the web pages to obtain the textual information related to the context in which the image appears on the Internet. The compiled textual information is often too long to be passed directly to the agent, and hence, we use the Llama-13B (Touvron et al. (2023)) language model to summarize this information. The summaries obtained from the LLM only focus on the most important parts of the text and hence also allow agents to develop a more focused understanding of the external context. While this method of summarization works for most samples in our dataset, there are some examples where the obtained web pages are not in the English language, and the LLM struggles with summarization. In this regard, we add an additional check that ignores text from web pages in languages other than English. While this restricts our system to the English language, it does not adversely affect system performance due to the distribution of the dataset, which consists of images mostly taken from English-language news articles. Multi-lingual support can be achieved by first translating text to English and then summarizing it.
@@ 3.4 Coherent Reasoning
All the different components of MAD-Sherlock are brought together in this stage of the pipeline. Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same. The agents also have access to external information related to the image through the external retrieval module. The final decision of the system is obtained once the debate terminates, which is after a certain number of debate rounds or after all agents converge to a common response, whichever is earlier.
@@ 4.1 Dataset
We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al. (2021)). The dataset is built based on the VisualNews (Liu et al. (2020)) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian and The Washington Post. The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair. CLIP (Radford et al. (2021)) is used to retrieve semantically similar images for a given caption. We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples. The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively.
@@ 4.2 Experimental Setup
All experiments were run on 8 A40 (46GB) Nvidia GPU server.
@@
We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al. (2024b)). The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
@@
We use the Bing Visual Search API (vis ) to run an image-based reverse search. Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new ). Finally, we use Llama-13B (Touvron et al. (2023)) to summarise the text obtained from the top $k=3$ web pages. This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
@@
We compare MAD-Sherlock to existing pretrained multi-modal baselines including CLIP (Radford et al. (2021)), VisualBERT (Li et al. (2019)), InstructBLIP (Dai et al. (2023)) and LLaVA (Liu et al. (2024b)). We also compare performance against GPT-4o (OpenAI & et al. (2024); OpenAI ). The models are presented with the image and caption pair and asked if the pair is misinformation. The models are further prompted to explain their reasoning. We also show results for two baseline methods trained from scratch, namely EANN (Wang et al. (2018)) and SAFE (Zhou et al. (2020)). We further compare MAD-Sherlock to DT-Transformer (Papadopoulos et al. (2023)), CCN (Abdelnabi et al. (2022)), Sniffer (Qi et al. (2024)), VINVL (Huang et al. (2024)), SSDL (Mu et al. (2023)) and Neuro-Sym (Zhu et al. (2022)).
@@ 4.3 Results
We present results for the experiments conducted to select the best debate setup as well compare the performance of MAD-Sherlock against existing methods. We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
@@ 4.3.1 Comparing Debate Setups
Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations. We also report results for the Asynchronous Debate setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain. The external retrieval of information significantly boosts performance. All following debate set-ups, therefore, use external information related to the image-caption pair as a part of their initial prompt. We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent. Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human. Further, the Asynchronous Debate setup benefits from the ensemble of agents which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses. The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought. We believe this accounts for the counter-intuitively performance of this method where agents perform worse with more information. The judged debate setup focuses on enforcing agents to structure their responses in a way that will convince a judge. The agents also debate with opposite stances and do not have the option of changing their stance mid-debate. This can further confuse the judge and lead to incorrect decisions. This is resolved in the Async_Debatehuman set up where agents are given complete freedom over their initial opinions, as well as their opinions during the debate. If they believe they are convinced by the other agents’ arguments, they can choose to change their response and the debate ends. Based on the results from Table 1, we choose the best-performing debate set-up, i.e. Async_Debatehuman (with external information) as the debate configuration for further experimentation and comparison.
@@ 4.3.2 Performance Comparison
MAD-Sherlock (ours) 90.17
Table 2: Performance comparison between our model and baselines: MAD-Sherlock (with GPT-4o) out performs all related work.
Table 2 shows the comparison between our system and existing methods. We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al. (2024); OpenAI ) model. Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset. While we do not provide a quantitative assessment of explanations by MAD-Sherlock, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines. We leave the detailed analysis of these explanations and the development of the associated metrics as future work.
We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work. This can be attributed to the necessity for external context for misinformation detection in the news domain and the lack of diverse perspectives that arise naturally in a multi-agent framework. Therefore, these standalone models, while promising, currently are unable to detect misinformation effectively. These models require additional integration into more comprehensive pipelines, as done in this work. In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al. (2020)) and EANN (Wang et al. (2018)) perform worse than pretrained multi-modal models. This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
@@ 5 User Study
We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation. While it is easy to quantify model performance in terms of misinformation detection, there are no effective metrics to assess the quality of the explanations generated by the model. Therefore, in order to perform a thorough analysis of the system performance, a user study is essential. For a deeper analysis we further grouped the participants based on their profession into three groups, namely: Journalists, AI Academics (studying AI) and Others. Further details regarding the study setup and participant groups can be found in Appendix A.4.
In the study, participants were shown ten image-text pairs and were asked to decide if the image and caption, when considered together, were misinformation or not. They were also asked to provide a confidence rating for their answer on a scale of 0-10, with 10 being the highest confidence level. For each image-text pair, after the participants provided their initial answers, they were shown AI insights about the same image-text pair. These AI insights were the final output explanations from MAD-Sherlock. Participants were then asked to reconsider their answers in light of the new information from the AI agent. Table 3 shows that average system performance is better than the average human performance for both cases where the participants have access to AI insights and where they do not. Therefore, MAD-Sherlock can be used as a reliable assistive tool for OSINT research for detecting and explaining misinformation with little or no human intervention.
Study Setup Average Accuracy
@@
Humans+MAD-Sherlock $76.7\pm 12.2$
MAD-Sherlock $\mathbf{80.0\pm 0.0}$
Table 3: Performance comparison between different study setups: MAD-Sherlock outperforms humans with and without AI assistance.
We further observe that the average human accuracy for the misinformation detection task increases by more than 27% concretizing the fact that AI insights from our model do actually improve human efficiency in detecting misinformation. We also observe interesting patterns for group-wise analysis which we believe would be valuable for future work. Table 4 shows that the performance of all groups improves significantly and is not far off from that of professional journalists. The average confidence level (out of 10) is comparable across all the groups before and after considering MAD-Sherlock insights and generally increases. Therefore, we conclude that MAD-Sherlock can significantly uplift non-expert performance and hence can be useful in citizen intelligence applications.
Avg acc $\uparrow$ Avg conf $\uparrow$ Avg acc $\uparrow$ Avg conf $\uparrow$
Group (only human) (only human) (with (with
MAD-Sherlock) MAD-Sherlock)
Journalists $70.0\pm 1.4$ $4.3\pm 2.1$ $82.2\pm 0.9$ $5.3\pm 1.3$
AI Academics $60.7\pm 1.4$ $3.2\pm 0.8$ $79.3\pm 1.3$ $5.8\pm 1.4$
Others $56.7\pm 1.5$ $3.9\pm 1.2$ $71.7\pm 1.1$ $5.8\pm 1.4$
Table 4: Performance comparison between different participant groups: All groups show performance improvement with MAD-Sherlock. AI Academics are able to perform nearly at par with professional journalists after considering insights from MAD-Sherlock.
@@ 6 Conclusion and Future Work
Our final system, MAD-Sherlock, achieves state-of-the-art performance on the misinformation detection task. Further, owing to our advanced external retrieval module, MAD-Sherlock provides clear, coherent and detailed explanations. As a result, MAD-Sherlock significantly improves the OOC misinformation detection performance of both human experts, and non-experts.
We identify several promising avenues for future research in this field. Notably, improving how agents handle disambiguation queries could significantly enhance detection accuracy and explainability by providing access to additional information. The research community would benefit from a continuously updated benchmark dataset, incorporating more recent news articles and subtler inconsistencies. A direct extension of this work involves applying our methods to video-text pairs. Since MAD-Sherlock’s explanations are currently limited to text, future research should explore multi-modal and mixed-modal explanation generation. Additionally, further experiments with different hyper-parameters, such as varying the number of agents and debate rounds, could refine the effectiveness of our approach.
Finally, while we conducted extensive user studies with MAD-Sherlock, deploying it on a larger scale in professional environments and within the citizen intelligence community will provide valuable insights into its real-world performance, uncovering new opportunities for improvement. For an analysis of limitations, please refer to Appendix A.1.
7 Acknowledgments
A part of this work was supported by UKRI DECaDE EP/T022485/1 and the UKRI grant: Turing AI Fellowship EP/W002981/1, Armasuisse Science+Technology. We would also like to thank the Royal Academy of Engineering and FiveAI.
@@ A.1 Limitations
Despite the strong performance of MAD-Sherlock, several limitations remain. First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve. Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration. Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
@@
Another important limitation is the potential risk that open-sourcing MAD-Sherlock might allow adversaries to train models specifically designed to counter or evade detection by our system. As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness. However, despite these risks, we believe that open-sourcing remains the right path forward. Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining MAD-Sherlock’s effectiveness in the long term. The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of MAD-Sherlock.

### v2

@@ 3.3.1 API-Based Information Retrieval
The Bing Visual Search API is used for the task of obtaining web pages related to a given image (vis, ) . A given image from the dataset is used to obtain a list of web pages completely and partially related to the image. We take the top three matching web pages in which the image appears. We believe these web pages contain sufficient information to allow the agent to develop a general understanding of the context in which the image is originally used. Since the community-accepted dataset for this task; NewsCLIPpings (Luo et al., 2021), contains images from articles published more than ten years ago, some images do not result in any web pages or viable search results. In such a scenario, we simply do not pass any external context to the agent and only rely on the agent’s existing knowledge base. Since this is relevant for only a very small number of examples in the dataset, it does not adversely affect system performance.
@@ 3.3.2 Summarization using LLM
Once the top three web pages are identified, we scrape their text to gather contextual information about the image. This textual data is often too lengthy for direct input to the agent, so we use the Llama-13B (Touvron et al., 2023) model to generate concise summaries, focusing on key details to help the agent better understand the external context. For non-English web pages, the LLM struggles with summarization, so we add a filter to exclude them. Although this restricts the system to English, it does not impact performance due to the dataset’s predominantly English-language content. Multi-lingual support could be added by translating text to English before summarization.
@@ 3.4 Coherent Reasoning
All the different components of LLM-Consensus are brought together in this stage of the pipeline. Each multi-modal agent is employed to participate in the best-debating set-up with the relevant prompts and is asked to detect a given image-text pair as misinformation and provide an explanation for the same. The agents also have access to external information related to the image through the external retrieval module. The final decision of the system is obtained once the debate terminates, which is after a certain number of debate rounds or after all agents converge to a common response, whichever is earlier.
@@ 4.1 Dataset
We perform a series of experiments and report results on the NewsCLIPpings dataset (Luo et al., 2021). The dataset is built based on the VisualNews (Liu et al., 2020) dataset, which consists of image-caption pairs from four news agencies: BBC, USA Today, The Guardian, and The Washington Post. The NewsCLIPpings dataset is created by generating OOC samples by replacing an image in one image-caption pair with a semantically related image from a different image-caption pair. CLIP (Radford et al., 2021) is used to retrieve semantically similar images for a given caption. We report results on the Merged-Balanced version of the dataset, which has balanced proportions of all the retrieval strategies and positive/negative samples. The training, validation and test sets have 71,072, 7,024 and 7,264 samples, respectively. We use the NewsCLIPpings dataset for evaluating our model’s efficacy in identifying out-of-context images and information due to its well-established use in prior work, its relative scale compared to other available datasets, and the lack of more recent large-scale benchmarks in this domain. Though no single dataset can capture the full diversity of real-world misinformation, NewsCLIPpings provides a standardized evaluation setting that facilitates comparability with existing methods.
@@ 4.2 Experimental Setup
All experiments were run on 8 A40 (46GB) Nvidia GPU server. The estimated cost of processing one data sample using LLM-Consensus is \$0.24. Inference times range from 5 to 15 seconds.
@@
We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al., 2024b). The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
@@
We use the Bing Visual Search API (vis, ) to run an image-based reverse search. Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new, ). Finally, we use Llama-13B (Touvron et al., 2023) to summarise the text obtained from the top $k=3$ web pages. This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
@@
We compare LLM-Consensus to existing pretrained multi-modal baselines including CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2024b). We also compare performance against GPT-4o (OpenAI & et al., 2024; OpenAI, ). The models are presented with the image and caption pair and asked if the pair is misinformation. The models are further prompted to explain their reasoning. We also show results for two baseline methods trained from scratch, namely EANN (Wang et al., 2018) and SAFE (Zhou et al., 2020). We further compare LLM-Consensus to DT-Transformer (Papadopoulos et al., 2023), CCN (Abdelnabi et al., 2022), Sniffer (Qi et al., 2024), VINVL (Huang et al., 2024), SSDL (Mu et al., 2023) and Neuro-Sym (Zhu et al., 2022).
@@ 4.3 Results
We present results for the experiments conducted to select the best debate setup as well compare the performance of LLM-Consensus against existing methods. We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
@@ 4.3.1 Comparing Debate Setups
Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations. We also report results for the Async_Debatehuman setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain. The external retrieval of information significantly boosts performance. We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent. Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human. Further, the asynchronous debate setup benefits from the ensemble of agents, which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses. The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought. We believe this accounts for the counter-intuitively performance of this method where agents perform worse with more information. The judged debate setup focuses on enforcing agents to structure their responses in a way that will convince a judge. The agents also debate with opposite stances and do not have the option of changing their stance mid-debate. This can further confuse the judge and lead to incorrect decisions. This is resolved in the Async_Debatehuman set up where agents are given complete freedom over their initial opinions, as well as their opinions during the debate. If they believe they are convinced by the other agents’ arguments, they can choose to change their response and the debate ends. Based on the results from Table 1, we choose the best-performing debate set-up, i.e. Async_Debatehuman (with external information) as the debate configuration for further experimentation and comparison.
@@ 4.3.2 Performance Comparison
LLM-Consensus (ours) 90.8
Table 2: Performance comparison between our model and baselines: LLM-Consensus (with GPT-4o) out performs all related work. Note: the GPT-4o# setup is identical to LLM-Consensus with the absence of a multi-agent debate, here only a single agent with access to external information is considered.
Table 2 shows the comparison between our system and existing methods. We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al., 2024; OpenAI, ) model. Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset. While we do not provide a quantitative assessment of explanations by LLM-Consensus, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines. This is attributed to the fact that in a multi-agent setup, we have multiple context windows which leads to more coherent and relevant final explanations. We leave the detailed analysis of these explanations and the development of the associated metrics as future work. We also note that the debate paradigm in itself is essential to the system performance. We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model.
We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work. This can be attributed to the necessity for external context for misinformation detection in the news domain and the lack of diverse perspectives that arise naturally in a multi-agent framework. Therefore, these standalone models, while promising, currently are unable to detect misinformation effectively. These models require additional integration into more comprehensive pipelines, as done in this work. In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al., 2020) and EANN (Wang et al., 2018) perform worse than pretrained multi-modal models. This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
4.3.3 Ablations
To analyze the importance of each component of the LLM-Consensus framework, we conduct ablation experiments. Specifically, we evaluate the effect of using LLaVA against GPT-4o, the impact of the external retrieval module, and the power of the multi-agent debate framework.
LLaVA GPT-4o Retrieval Debate Accuracy Precision Recall
$77.1$ $68.4$ $89.3$
$86.2$ $82.6$ $90.6$
$86.0$ $80.2$ $95.6$
$90.2$ $\mathbf{90.3}$ $90.1$
$\mathbf{90.8}$ $85.5$ $\mathbf{99.0}$
Table 3: Ablation: Quantitative evaluation of each component of LLM-Consensus on classification performance.
We observe that the combination of GPT-4o, the external retrieval module, and the multi-agent debate framework yields the highest performance across all metrics, with 90.8% accuracy, 85.5% precision, and 99.0% recall, demonstrating the value of combining these components. The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7 (as seen in Table 2) to 90.2, underscoring its role in enabling contextual reasoning and refining predictions. Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3 to 85.5 and recall from 90.1 to 99.0. Meanwhile, retrieval contributes more substantially to LLaVA’s performance gains, likely due to GPT-4o’s broader world knowledge. Without external retrieval or the debate framework, the performance drops, emphasizing the critical role of these components in achieving state-of-the-art results.
@@ 5 User Study
We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation. While model performance in misinformation detection is easy to quantify, there are no established metrics for assessing the quality of its explanations, making the user study essential. Participants were grouped into three categories based on their professions: Journalists, AI Academics, and Others (see Appendix A.4 for details).
During the study, participants reviewed ten image-text pairs and decided whether each pair constituted misinformation. They also rated their confidence on a scale of 0 to 10. After submitting their initial answers, participants were shown LLM-Consensus’s AI-generated explanations and asked to reconsider their responses. As shown in Table 4, the system outperformed the average human performance in both cases, with and without AI insights. These results demonstrate LLM-Consensus’s potential as a reliable assistive tool for OSINT research, capable of detecting and explaining misinformation with minimal human intervention.
[…]
@@
Humans+LLM-Consensus $76.7\pm 12.2$
LLM-Consensus $\mathbf{80.0\pm 0.0}$
Table 4: Performance comparison between different study setups: LLM-Consensus outperforms humans with and without AI assistance.
Group-wise analysis, shown in Table 5, reveals significant performance gains across all groups, with results approaching those of professional journalists. Confidence levels (out of 10) are comparable across groups and generally increase after using LLM-Consensus insights. Thus, LLM-Consensus can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
Metric Journalists Academics Others
Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
Accuracy (with LLM-Consensus) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
Confidence (with LLM-Consensus) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
Table 5: Performance comparison: LLM-Consensus improves performance across all participant groups.
@@ 6 Conclusion and Future Work
Our final system, LLM-Consensus, achieves state-of-the-art performance in misinformation detection while offering clear, detailed explanations, thanks to our advanced external retrieval module. As a result, we observe significant improvements in out-of-context (OOC) detection for both experts and non-experts.
We identify several promising directions for future research. The community would benefit from a continuously updated benchmark dataset incorporating recent news articles and subtler inconsistencies. We plan to extend our methods to video-text pairs and multi-lingual content. We also believe that integrating more advanced models in the summarization pipeline will further enhance performance and are interested in comparing LLM-Consensus with multi-agent systems using external retrieval.
Finally, we conducted extensive user studies but recognize the importance of large-scale deployment in professional settings and the citizen intelligence community. This will provide valuable real-world feedback and uncover further opportunities for improvement. For a discussion of limitations, see Appendix A.1.
@@ A.1 Limitations
Despite the strong performance of LLM-Consensus, several limitations remain. First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve. Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration. Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
@@
Another important limitation is the potential risk that open-sourcing LLM-Consensus might allow adversaries to train models specifically designed to counter or evade detection by our system. As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness. However, despite these risks, we believe that open-sourcing remains the right path forward. Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining LLM-Consensus’s effectiveness in the long term. The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of LLM-Consensus.